Status Report

ISS On-Orbit Status 31 Jul 2002

By SpaceRef Editor
July 31, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously.

Before crew breakfast, FE-2 Sergei Treschev started out the day by again activating the turbopump of the Russian Plasma-Crystal 3 crystallization experiment.  Later, he set up experiment operations and ran the experiment.  [The study consists of three major segments: ‘heart-beat’ instability in the dust cloud at various plasma discharge energies; study of residual charge attached to small particles after termination of discharge (relaxation); and repeat of segment 2, but with large particles.]  During the day, Sergei monitored chamber pressure using the VK-316M vacuum gauges.  After experiment completion, he tested and calibrated the payload, copied the data to the computer and removed the videocassette that was used to record the oscillating plasma cloud.  Before dinner, the turbopump was turned off again.

CDR Valery Korzun worked for several hours on the Russian Elektron oxygen generator system.  [After first preparing the unit for subsystem removal and replacement, he detached it from the BITS2-12 onboard telemetry system by demating its TMI-connector.  Assisted by FE-2 Treschev, Korzun then removed the BZh Fluids Unit #4 and replaced it with Fluids Unit #3, tagging up with ground specialists via S-band. BITS2-12 was then reconnected to the system.  Later in the day, the new Fluids Unit was to be pressurized and Elektron activated from laptop for testing, with its electrolytic cell voltage being monitored every 10 minutes for 1.5 hours.]

SUBSA (solidification using a baffle in sealed ampoules) troubleshooting went very well yesterday.  With Peggy Whitson’s help, the ground confirmed that a local temperature setpoint limit in the Eurotherm PCM (process control module) prevented the SUBSA TC (thermal chamber) from heating up above 825C.  The setpoint limit was successfully reconfigured to 850C.  Today, the ground ran a SUBSA Eurotherm controller verification test to make sure that the troubleshooting worked and that the TC can be commanded to a temperature of 845C.  Peggy was asked not to insert a sample into the TC for this exercise.  An abbreviated operations procedure was used to direct her through the power up cycle in the correct sequence.

In preparation for return of a box with two PPO InPU flash cards, CDR Korzun was requested to look for them and confirm their stowage location.  The flash cards are currently planned to be returned on Soyuz-208.

FE-1 Whitson concluded this week’s water sampling by analyzing the MCD (microbial capture device) samples for their T+2 days reading.  Her microbiological analysis data were then transferred to the MEC (medical equipment computer).  T+5d analysis is not required.

She also did routine maintenance on the CSA-CP (compound specific analyzer-combustion products).

In preparation for tomorrow’s planned SSRMS (space station remote manipulator system) operations, CDR Korzun and FE-1 Whitson had an hour on today’s schedule for studying the applicable new DOUG (dynamic operations ubiquitous graphics) setup.  [Tomorrow, after the crew has rebooted the Lab and CUPola RWS (robotic workstation) laptops and the ground has powered up the SSRMS, operations are planned to complete several open OCRs (On-Orbit Checkout Requirements) for the MBS/SSRMS (mobile base system/SSRMS).  The crew will release MBS PDGF 4 (power and data grapple fixture #4), followed by a series of maneuvers, calibrations and checkouts.  Finally, PDGRF 2 will be grappled and released.]

FE-1 Whitson set up the ARIS ICE (active rack isolation system/isolation characterization experiment) after the ground powered down the ARIS controller.    Later in the day, she removed the ARIS alignment guides. Tomorrow morning, POC (Payload Operations Center) is expected to confirm that ARIS is in a good configuration for Peggy to proceed.

CDR Korzun and FE-2 Treschev completed their weekly NTXN "Interactions" questionnaires, while FE-1 Whitson filled in her periodic FFQ (food frequency questionnaire) that keeps a running account of her nutritional intake.

At 10:10 am EDT today, Peggy Whitson will downlink a five-minute greeting to the Houston Astros, followed by a 25-min. TV interview with station  KMEG-TV in Sioux City, Iowa, using KU-band/video and S-band/audio.

Daily routine servicing tasks today were performed by Korzun (SOSh life support system maintenance), Treschev (IMS delta file preparation) and Whitson (Lab payload status checkup).

All crewmembers did their physical exercise program on RED, VELO and CEVIS (cycle ergometer with vibration isolation).

TVIS troubleshooting:  Still no good news.  MCC-H analysis showed that the truss could fracture within hours of use. All assumptions and input into that analysis are being reviewed, but even the best case only gives a few weeks of use.  Primary concern is to develop a fix that will not break causing irreparable damage to the TVIS or, more significantly, harm to the runner. This concern is pushing engineers away from doing anything with TVIS prior to the next launch (9A or Progress 9P).  Current focus is to design a fix to be brought up on the next available launch, to ensure that there is an operable TVIS on board for the end of the current crew’s stay, rather than work up a quick fix that fails and precludes TVIS ops altogether during their your stay.  However, workable and safe near term solutions are still being sought.

SpaceRef staff editor.